OpenStudy (anonymous):

Simplify the expression: cot x sin x - sin quantity pi divided by two minus x. + cos x cos x sin x 2 sin x 2 cos x

OpenStudy (anonymous):

Turn cot(x) into sines and cosines and then know that: \[\sin(\frac{ \pi }{ 2 }- \theta ) = \cos \theta\]

OpenStudy (dumbcow):

and cot = 1/tan = cos/sin